The ages of shelby and her father combined equals 55 years. Shelby's father's age,f, is one less than three times Shelby's age,s, What are their ages?
could you please explain how to do this

Answer by mananth(16949)

Shelby's age = x
shelby's father age = 55-x
...
55-x= 3x-1
3x+x=55+1
4x=56
/4
x=14 shelby's age
father's age = 55-14 = 41 years
m.ananth@hotmail.ca